A spinner is divided into 8 congruent regions numbered 1 through 8. The spinner is spun once.Find the probability of each event.
1. The number is even or divisible by 3.

Answer by stanbon(48568)

A spinner is divided into 8 congruent regions numbered 1 through 8. The spinner is spun once.Find the probability of each event.
1. The number is even or divisible by 3.
Evens are sections 2,4,6,8
Divisible by 3 are 3,6
Prob(even or divisible by 3)=Prob(even)+Prob(divis by 3)-Prob(even and divis by 3)
=4/8+2/8-1/8=5/8
Cheers,
Stan H.